OpenStudy (anonymous):

Is my answer correct? :) Solve. 2x2 - x = 15 Answer: x = 3 and x = 1

jimthompson5910 (jim_thompson5910):

that's incorrect

jimthompson5910 (jim_thompson5910):

if you plug x = 1 into the equation, you get 2x^2-x = 15 2(1)^2-1 = 15 2(1)-1 = 15 2 - 1 = 15 1 = 15 which is FALSE so x = 1 is NOT a solution

jimthompson5910 (jim_thompson5910):

However, x = 3 is a solution because.... 2x^2-x = 15 2(3)^2-3 = 15 2(9)-3 = 15 18 - 3 = 15 15 = 15 which is TRUE so x = 3 is a solution

jimthompson5910 (jim_thompson5910):

So you can use this idea to test possible solutions

OpenStudy (anonymous):

x= 7/2 x = -1 ?

jimthompson5910 (jim_thompson5910):

2x^2-x = 15 only has 2 solutions and I just pointed one of those solutions out to you

OpenStudy (anonymous):

x= -5/2

OpenStudy (anonymous):

x=3

jimthompson5910 (jim_thompson5910):

perfect, you got it